The wall of an oven operating at 1000°C consists of two layers of different material. The inside layer of wall is built of 40cm of kaoline firebrick having k-0.08 W/m-K, surrounded by 40cm of fireclay brick having k-0,8 W/m-K. The outer side of the wall is maintained at Process Heat Transfer 50°C. What is the temperature at the interface of the layers?_____


1
Expert's answer
2021-08-25T15:31:01-0400

"R_{th1} =\dfrac{l_1}{k_1A}=\dfrac{0.40}{0.8} = 0.5\ K/W"


"Q =\dfrac{T_1-T_0}{R_{th1}} =\dfrac{1000-50}{0.5} = 1900W"


"T_1-T_0 \textsf{ (interface)}= 950°C"